# Jointly Computation- and Communication-Efficient Distributed Learning

### v1

@@ I Introduction
Smart devices equipped with computational and communications resources underwent, in recent years, a widespread adoption in many applications, including power grids, robotics, traffic and sensor networks [1, 2]. These devices then become the components of multi-agent systems that can collect data and cooperatively achieve learning tasks. However, their resources (CPU and communications) might be limited. Thus, in this paper we focus on the design of a distributed learning algorithm that is jointly computation- and communication-efficient.
Different techniques have been explored to design computation-efficient algorithms. The main solution is the use of stochastic gradients, which allow the agents to update their models using only a subset of their local data [3, 4]. However, stochastic gradients might cause inexact convergence, and variance reduction was proposed to solve this issue, see e.g., [5, 6].
In terms of communication-efficiency, the main methods adopted in machine learning are compression and local training. Compression allows to reduce the size of communications exchanged by the agents, for example by transmitting only the most significant components of a model. Different distributed algorithms were designed to use compressed communications (see, for instance [7, 8, 9, 10]. Of these, [7, 8] also implement an error feedback mechanism, which ensures exact convergence even when using compression.
While compression reduces the size of communications, local training instead reduces their frequency. The idea is to allow the agents to perform multiple steps of training between each round of communications. This paradigm has been applied e.g. in [11, 12] to gradient tracking, in [4] to the distributed dual ascent method, and in our works [13, 14] that focus on redesigning the distributed ADMM.
@@
  • We propose a novel algorithm, LT-ADMM-CC (Local Training ADMM with Compressed Communication), based on  [14], which employs stochastic gradients with variance reduction for computation-efficiency, and compression and local training for communication-efficiency. Additionally, our design integrates error feedback.
@@ II-A Problem Formulation
The cost $f_{i}$ of each agent $i\in\mathcal{V}$ is $L$-smooth and $\mu$-strongly convex, with $0<\mu\leq L<\infty$.111 We recall that a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is $L$-smooth if it is differentiable and $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|$, $\forall x,y\in\mathbb{R}^{n}$; moreover, $f$ is $\mu$-strongly convex if $\left\langle\nabla f_{i}(x)-\nabla f_{i}(y),x-y\right\rangle\geq\mu\|x-y\|^{2}$, $\forall x,y\in\mathbb{R}^{n}$.
@@
These assumptions are standard for distributed problems, with only strong convexity being somewhat restrictive. However, strong convexity is instrumental to prove linear convergence; and we remark that it can be relaxed to allow for nonconvex problems extending the analysis of [14].
@@ II-B Algorithm design
We recall that we focus on designing a distributed learning algorithm that is jointly computation- and communication-efficient. To this end, we start from the distributed ADMM of [15], and in the following steps we integrate some suitable design modifications. Therefore, our starting point is the algorithm characterized by the updates:
@@
The distributed implementation of (3) entails agent $j\in\mathcal{N}_{i}$ sending $z_{ji,k}+2\rho x_{j,k+1}$ to $i$ [15]. However, in a learning setting where $n\gg 1$, this requires prohibitively large bandwidth, and as a solution compressed communications can be employed. Letting $\mathcal{C}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a compression operator, [^2] Precise assumptions will be introduced in section II-C. we could then allow agent $j$ to transmit $\mathcal{C}\left(-z_{ji,k}+2\rho x_{j,k+1}\right)$. This design choice would reduce the communication burden, but would also result in inexact convergence. Thus, we also integrate an error feedback mechanism that asymptotically reduces to zero the error induced by compression. In particular, we define the auxiliary variables $u_{i,k}$ and $s_{ij,k}$, and rewrite (3b) as:
@@
Besides reducing the required bandwidth via compression, in our design we also reduce the frequency of communications via local training. Following the idea in [14], we notice that update (3a) requires the solution of an optimization problem, which in general lacks a closed form. The idea then is to allow the agents to approximate its solution with a finite number $\tau$ of gradient-based steps:
@@
The use of full gradient evaluations $g_{i}(\phi_{i,k}^{t})=\nabla f_{i}(\phi_{i,k}^{t})$ might be computationally expensive when $m_{i}\gg 1$ in (1). Thus, we allow the agents to use a variance reduced stochastic gradient estimator [16], where each agent maintains a table of component gradients $\{\nabla f_{i,h}(r_{i,h,k}^{t})\},h=1,\ldots,m_{i}$, where $r_{i,h,k}^{t}$ represents the most recent iterate at which the component gradient was computed. This table is reset at the start of each new local training, and the agents estimate their local gradients as
@@ II-C Convergence analysis
The compression operator $\mathcal{C}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is unbiased $\mathbb{E}\bigl{[}\mathcal{C}(x)\bigr{]}=x$, and there exists a constant $p\geq 1$ such that $\mathbb{E}\bigl{[}\|\mathcal{C}(x)-x\|^{2}\bigr{]}\leq p\|x\|^{2},\forall x\in\mathbb{R}^{n}.$
@@ III-B Comparison
We compare now the performance of LT-ADMM-CC with alternative distributed optimization methods that employ compression: CEDAS [9], COLD [8], DPDC [7, Algorithm 1], and LEAD [10]. For all algorithms, we use an $8$-bit quantizer and stochastic gradients with batch-size $|\mathcal{B}|=1$; we also hand-tune the parameters of the algorithms to achieve optimal performance.
@@
LEAD [10] $\tau(t_{g}+t_{c})$
CEDAS [9] $\tau(t_{g}+2t_{c})$
COLD [8]& DPDC [7] $\tau\left(t_{g}+t_{c}\right)$ stochastic gradient
COLD [8]& DPDC [7] $\tau\left(m_{i}t_{g}+t_{c}\right)$ full gradient
@@ IV Concluding Remarks
Refer to [14, Lemma 1], we know that if a matrix $\widehat{\mathbf{Q}}\in\mathbf{R}^{N\times(N-1)}$ the matrix satisfying $\widehat{\mathbf{Q}}\widehat{\mathbf{Q}}^{T}=\mathbf{I}_{N}-\frac{1}{N}\mathbf{11}{}^{T}$, $\widehat{\mathbf{Q}}^{T}\widehat{\mathbf{Q}}=\mathbf{I}_{N-1}$ and $\mathbf{1}^{T}\widehat{\mathbf{Q}}=0$, $\widehat{\mathbf{Q}}^{T}\mathbf{1}=0$. We diagonalize $\mathbf{D}_{i}=\mathbf{V}_{i}\mathbf{\Delta}_{i}\mathbf{V}_{i}^{-1}$, where $\mathbf{\Delta}_{i}$ is the diagonal matrix of $\mathbf{D}_{i}$’s eigenvalues, and
@@
with $d_{12}=-\beta\tau+((\beta\tilde{\lambda}_{i}\rho\tau(\beta\tilde{\lambda}_{i}\rho\tau+2))^{0.5})/(\tilde{\lambda}_{i}\rho)$, $d_{13}=-\beta\tau-((\beta\tilde{\lambda}_{i}\rho\tau(\beta\tilde{\lambda}_{i}\rho\tau+2))^{0.5})/(\tilde{\lambda}_{i}\rho)$, $d_{22}=\tilde{\lambda}_{i}\rho d_{12}-1$, $d_{23}=\tilde{\lambda}_{i}\rho d_{13}-1$, $\tilde{\lambda}_{i}<0$ is the nonzero eigenvalue of $\tilde{\mathbf{L}}$. We conclude that we can write $\mathbf{\Theta}=(\mathbf{P}_{0}\bm{\phi})^{T}\mathbf{V}\mathbf{\Delta}\mathbf{V}^{-1}(\mathbf{P}_{0}\bm{\phi})$ where $\mathbf{V}=\operatorname{blkdiag}\left\{V_{i}\right\}_{i=2}^{N}$ and $\mathbf{\Delta}=\operatorname{blkdiag}\left\{\mathbf{\Delta}_{i}\right\}_{i=2}^{N}$. Moreover, $\|\mathbf{\Delta}\|=1-{r^{2}\lambda_{l}\rho\tau\beta}/{2}$ when $r^{2}\lambda_{u}\rho\tau\beta<2$. $\widehat{\mathbf{V}}^{-1}=\mathbf{V}^{-1}(\mathbf{P}_{0}\bm{\phi})$, where $\mathbf{P}_{0}$ is a permutation matrix and $\bm{\phi}$ is an orthogonal matrix, yields $\widehat{\mathbf{d}}_{k+1}=\mathbf{\Delta}\widehat{\mathbf{d}}_{k}-\widehat{\mathbf{h}}_{k},$ where $\widehat{\mathbf{d}}_{k}=\widehat{\mathbf{V}}^{-1}\begin{bmatrix}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k};\widehat{\mathbf{Q}}^{T}\mathbf{Y}_{k};\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{Y}}_{k}\end{bmatrix}$, $\widehat{\mathbf{h}}_{k}=\widehat{\mathbf{V}}^{-1}\widehat{\mathbf{Q}}^{T}\mathbf{h}_{k}$. ∎
@@
Using similar analysis with [4, Lemma 2] and (21), we obtain
@@
Since $\langle(z-y),\nabla g(x)\rangle\geq g(z)-g(y)+\frac{\mu}{4}\|y-z\|^{2}-L\|z-x\|^{2},\forall x,y,z\in R^{n}$ for any L-smooth and $\mu$-strongly convex function $g$ [17], we have

### v2

@@ I Introduction
Smart devices equipped with computational and communications resources underwent, in recent years, a widespread adoption in many applications, including power grids, robotics, traffic and sensor networks [molzahn_survey_2017, nedic_distributed_2018]. These devices then become the components of multi-agent systems that can collect data and cooperatively achieve learning tasks. However, their resources (CPU and communications) might be limited. Thus, in this paper we focus on the design of a distributed learning algorithm that is jointly computation- and communication-efficient.
Different techniques have been explored to design computation-efficient algorithms. The main solution is the use of stochastic gradients, which allow the agents to update their models using only a subset of their local data [koloskova2019decentralized, alghunaim_local_2023]. However, stochastic gradients might cause inexact convergence, and variance reduction was proposed to solve this issue, see e.g., [li_variance_2022, jiang_distributed_2023].
In terms of communication-efficiency, the main methods adopted in machine learning are compression and local training. Compression allows to reduce the size of communications exchanged by the agents, for example by transmitting only the most significant components of a model. Different distributed algorithms were designed to use compressed communications (see, for instance [yi2022communication, zhang2023innovation, huang2024cedas, liulinear]. Of these, [yi2022communication, zhang2023innovation] also implement an error feedback mechanism, which ensures exact convergence even when using compression.
While compression reduces the size of communications, local training instead reduces their frequency. The idea is to allow the agents to perform multiple steps of training between each round of communications. This paradigm has been applied e.g. in [hien_nguyen_performance_2023, liu_decentralized_2023] to gradient tracking, in [alghunaim_local_2023] to the distributed dual ascent method, and in our works [ren2024distributed, ren2025communication] that focus on redesigning the distributed ADMM.
@@
  • We propose a novel algorithm, LT-ADMM-CC (Local Training ADMM with Compressed Communication), based on  [ren2025communication], which employs stochastic gradients with variance reduction for computation-efficiency, and compression and local training for communication-efficiency. Additionally, our design integrates error feedback.
@@ II-A Problem Formulation
The cost $f_{i}$ of each agent $i\in\mathcal{V}$ is $L$-smooth and $\mu$-strongly convex, with $0<\mu\leq L<\infty$. [^1] We recall that a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is $L$-smooth if it is differentiable and $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|$, $\forall x,y\in\mathbb{R}^{n}$; moreover, $f$ is $\mu$-strongly convex if $\left\langle\nabla f_{i}(x)-\nabla f_{i}(y),x-y\right\rangle\geq\mu\|x-y\|^{2}$, $\forall x,y\in\mathbb{R}^{n}$.
@@
These assumptions are standard for distributed problems, with only strong convexity being somewhat restrictive. However, strong convexity is instrumental to prove linear convergence; and we remark that it can be relaxed to allow for nonconvex problems extending the analysis of [ren2025communication].
@@ II-B Algorithm design
We recall that we focus on designing a distributed learning algorithm that is jointly computation- and communication-efficient. To this end, we start from the distributed ADMM of [bastianello_asynchronous_2021], and in the following steps we integrate some suitable design modifications. Therefore, our starting point is the algorithm characterized by the updates:
@@
The distributed implementation of (3) entails agent $j\in\mathcal{N}_{i}$ sending $z_{ji,k}+2\rho x_{j,k+1}$ to $i$ [bastianello_asynchronous_2021]. However, in a learning setting where $n\gg 1$, this requires prohibitively large bandwidth, and as a solution compressed communications can be employed. Letting $\mathcal{C}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a compression operator, [^2] Precise assumptions will be introduced in section II-C. we could then allow agent $j$ to transmit $\mathcal{C}\left(-z_{ji,k}+2\rho x_{j,k+1}\right)$. This design choice would reduce the communication burden, but would also result in inexact convergence. Thus, we also integrate an error feedback mechanism that asymptotically reduces to zero the error induced by compression. In particular, we define the auxiliary variables $u_{i,k}$ and $s_{ij,k}$, and rewrite (3b) as:
@@
Besides reducing the required bandwidth via compression, in our design we also reduce the frequency of communications via local training. Following the idea in [ren2025communication], we notice that update (3a) requires the solution of an optimization problem, which in general lacks a closed form. The idea then is to allow the agents to approximate its solution with a finite number $\tau$ of gradient-based steps:
@@
The use of full gradient evaluations $g_{i}(\phi_{i,k}^{t})=\nabla f_{i}(\phi_{i,k}^{t})$ might be computationally expensive when $m_{i}\gg 1$ in (1). Thus, we allow the agents to use a variance reduced stochastic gradient estimator [defazio2014saga], where each agent maintains a table of component gradients $\{\nabla f_{i,h}(r_{i,h,k}^{t})\},h=1,\ldots,m_{i}$, where $r_{i,h,k}^{t}$ represents the most recent iterate at which the component gradient was computed. This table is reset at the start of each new local training, and the agents estimate their local gradients as
@@ II-C Convergence analysis
The compression operator $\mathcal{C}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is unbiased $\mathbb{E}\bigl[\mathcal{C}(x)\bigr]=x$, and there exists a constant $p>1$ such that $\mathbb{E}\bigl[\|\mathcal{C}(x)\|^{2}\bigr]\leq p\|x\|^{2},\forall x\in\mathbb{R}^{n}.$
@@ III-B Comparison
We compare now the performance of LT-ADMM-CC with alternative distributed optimization methods that employ compression: CEDAS [huang2024cedas], COLD [zhang2023innovation], DPDC [yi2022communication, Algorithm 1], and LEAD [liulinear]. For all algorithms, we use an $8$-bit quantizer and stochastic gradients with batch-size $|\mathcal{B}|=1$; we also hand-tune the parameters of the algorithms to achieve optimal performance.
@@
LEAD [liulinear] $\tau(t_{g}+t_{c})$
CEDAS [huang2024cedas] $\tau(t_{g}+2t_{c})$
COLD [zhang2023innovation]& DPDC [yi2022communication] $\tau\left(t_{g}+t_{c}\right)$ stochastic gradient
COLD [zhang2023innovation]& DPDC [yi2022communication] $\tau\left(m_{i}t_{g}+t_{c}\right)$ full gradient
@@ IV Concluding Remarks
Refer to [ren2025communication, Lemma 1], we know that if a matrix $\widehat{\mathbf{Q}}\in\mathbf{R}^{N\times(N-1)}$ the matrix satisfying $\widehat{\mathbf{Q}}\widehat{\mathbf{Q}}^{T}=\mathbf{I}_{N}-\frac{1}{N}\mathbf{11}{}^{T}$, $\widehat{\mathbf{Q}}^{T}\widehat{\mathbf{Q}}=\mathbf{I}_{N-1}$ and $\mathbf{1}^{T}\widehat{\mathbf{Q}}=0$, $\widehat{\mathbf{Q}}^{T}\mathbf{1}=0$. We diagonalize $\mathbf{D}_{i}=\mathbf{V}_{i}\mathbf{\Delta}_{i}\mathbf{V}_{i}^{-1}$, where $\mathbf{\Delta}_{i}$ is the diagonal matrix of $\mathbf{D}_{i}$’s eigenvalues, and
@@
with $d_{12}=-\beta\tau+((\beta\tilde{\lambda}_{i}\rho\tau(\beta\tilde{\lambda}_{i}\rho\tau+2))^{0.5})/(\tilde{\lambda}_{i}\rho)$, $d_{13}=-\beta\tau-((\beta\tilde{\lambda}_{i}\rho\tau(\beta\tilde{\lambda}_{i}\rho\tau+2))^{0.5})/(\tilde{\lambda}_{i}\rho)$, $d_{22}=\tilde{\lambda}_{i}\rho d_{12}-1$, $d_{23}=\tilde{\lambda}_{i}\rho d_{13}-1$, $\tilde{\lambda}_{i}<0$ is the nonzero eigenvalue of $\tilde{\mathbf{L}}$. We conclude that we can write $\mathbf{\Theta}=(\mathbf{P}_{0}\bm{\phi})^{T}\mathbf{V}\bm{\Delta}\mathbf{V}^{-1}(\mathbf{P}_{0}\bm{\phi})$ where $\mathbf{V}=\operatorname{blkdiag}\left\{V_{i}\right\}_{i=2}^{N}$ and $\bm{\Delta}=\operatorname{blkdiag}\left\{\mathbf{\Delta}_{i}\right\}_{i=2}^{N}$. Moreover, $\|\mathbf{\Delta}\|=1-{r^{2}\lambda_{l}\rho\tau\beta}/{2}$ when $r^{2}\lambda_{u}\rho\tau\beta<2$. $\widehat{\mathbf{V}}^{-1}=\mathbf{V}^{-1}(\mathbf{P}_{0}\bm{\phi})$, where $\mathbf{P}_{0}$ is a permutation matrix and $\bm{\phi}$ is an orthogonal matrix, yields $\widehat{\mathbf{d}}_{k+1}=\mathbf{\Delta}\widehat{\mathbf{d}}_{k}-\widehat{\mathbf{h}}_{k},$ where $\widehat{\mathbf{d}}_{k}=\widehat{\mathbf{V}}^{-1}\begin{bmatrix}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k};\widehat{\mathbf{Q}}^{T}\mathbf{Y}_{k};\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{Y}}_{k}\end{bmatrix}$, $\widehat{\mathbf{h}}_{k}=\widehat{\mathbf{V}}^{-1}\widehat{\mathbf{Q}}^{T}\mathbf{h}_{k}$. ∎
@@
Using similar analysis with [alghunaim_local_2023, Lemma 2] and (21), we obtain
@@
Since $\langle(z-y),\nabla g(x)\rangle\geq g(z)-g(y)+\frac{\mu}{4}\|y-z\|^{2}-L\|z-x\|^{2},\forall x,y,z\in R^{n}$ for any L-smooth and $\mu$-strongly convex function $g$ [nesterov2013introductory], we have